# Color it, Code it, Cancel it: $k$-local dynamical decoupling from classical additive codes

### v1

@@ IV.1 Orthogonal arrays and dynamical decoupling
Time-efficient DD sequences can be constructed using combinatorial design objects. For example, Ref. [27] employs Hadamard matrices to construct decoupling sequences for two-local Hamiltonians. We generalize this approach to arbitrary $k$-local Hamiltonians by using more general combinatorial structures known as orthogonal arrays [30, 32, 43]. In the following analysis, we consider the worst-case scenario where all possible $k$-local interactions between the colored qubits $c_{i}$ are present in the quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$.
@@
An orthogonal array $\text{OA}(L,\chi_{I},S,k)$ is an $L\times\chi_{I}$ array with entries taken from a finite set $S$ and strength $k$, such that every possible $k$-tuple from $S^{\otimes k}$ appears uniformly across any choice of $k$ columns [44]. In this work, we restrict ourselves to the set $S=\{I,X,Y,Z\}$ of single-qubit Pauli operators. Each column of the orthogonal array is associated with a colored qubit $c_{i}$, and each row defines a group element in the DD group $\mathcal{G}$. If the noise model is biased along a specific axis of the Bloch sphere (e.g. the $Z$-axis for pure dephasing) or if we want to maintain certain interactions (e.g. $XX$ interactions), we can reduce the required size $L$ of the orthogonal array by selecting a smaller set (e.g. $S=\{I,X\}$) instead of the full single-qubit Pauli group. We remark that for any choice of parameters $\chi_{I}$, $S$, and strength $k$, there always exists an orthogonal array whose length $L$ is determined by these parameters [44].
@@
In Fig. 4, we show the explicit orthogonal array $\text{OA}(L=16,\chi_{I}=3,|S|=4,k=2)$, which generates a universal decoupling sequence for two-local Hamiltonians defined on the quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ [45]. One can verify that for any pair of columns, the rows of the OA exhaustively span all combinations of $\{I,X,Y,Z\}^{\otimes 2}$ exactly once. Moreover, each Pauli operator appears exactly four times in every column, ensuring balanced coverage and uniformity.
@@
By construction, the action of the DD group $\mathcal{G}$ derived from an OA with strength $k$ implements a group twirl over the $k$-fold Pauli group $\{I,X,Y,Z\}^{\otimes k}$ on any subset of $k$ colored qubits. As a consequence, all $k$-local interaction terms are dynamically canceled, following directly from Schur’s lemma [Eq. (10)]. Furthermore, all interactions with support on less than $k$ colored qubits are also canceled as a consequence of this construction. Therefore, to universally decouple all $k$-local interactions on the quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$, it suffices to construct the orthogonal array $\text{OA}(L,\chi_{I},S,k)$.
@@
We note that there is a comprehensive online library cataloging efficient orthogonal arrays across a wide range of parameters [45]. In the next section, we describe how to construct length-efficient orthogonal arrays by leveraging their connection to classical error-correcting codes. Since classical codes are extensively studied, we can utilize well-established classical codes to obtain time-efficient DD sequences suppressing arbitrary $k$-local interactions [30].
@@ IV.2 Orthogonal array and classical additive codes
The connection between orthogonal arrays and classical linear codes is well established [46, 44]. In this work, we focus on the single-qubit Pauli group modulo global phases, which admits a natural isomorphism to the finite field $\mathbb{F}_{4}$. A finite field $\mathbb{F}_{q}$ is a set containing $q$ elements in which addition, subtraction, multiplication, and division (except by zero) are all well-defined and satisfy the field axioms. The isomorphism between Pauli group (modulo phase) and $\mathbb{F}_{4}$ is defined by the correspondence
@@
This correspondence allows us to represent Pauli operators algebraically as elements of $\mathbb{F}_{4}$, providing a convenient framework for constructing orthogonal arrays from quaternary codes. Furthermore, when the noise is biased (e.g., dephasing along the $Z$-axis), we may restrict to a subset of the Pauli group, in which case classical codes over the binary field $\mathbb{F}_{2}$ become relevant.
We consider here classical additive codes, which, as we will show in Section IV.3, provide a natural framework for interpreting DD as an error detection code. An additive code $\mathcal{C}$ is a subset of $\mathbb{F}_{q}^{n}$ (a vector of length $n$ in which its entries are from $\mathbb{F}_{q}$) that is closed under addition. While every linear code is an additive code, the converse is not true: additive codes are not necessarily closed under scalar multiplication unless the field is $\mathbb{F}_{2}$.
@@
The vector space $\mathbb{F}_{4}^{n}$ can be endowed with an inner product $\langle\cdot,\cdot\rangle$. A natural choice for additive codes is the trace Hermitian inner product [47], defined as
$\langle\bf{u},\bf{v}\rangle_{tr}=\sum_{i}(u_{i}v_{i}^{2}+v_{i}u_{i}^{2}).$ (23)
Under the mapping between single-qubit Pauli operators and elements of $\mathbb{F}_{4}$ given in Eq. (19), this inner product directly encodes the commutation relations of Pauli operators, that is, it evaluates to zero if the corresponding Pauli operators commute and to one if they anticommute.
@@
$\mathcal{C}^{\perp}=\{\bf{v}\in\mathbb{F}_{q}^{n}~|~\langle\bf{u},\bf{v}\rangle_{tr}=0,\quad\forall\bf{u}\in\mathcal{C}\}.$ (24)
@@
A seminal result by Delsarte [46] establishes a fundamental connection between orthogonal arrays and classical codes:
@@
In Ref. [30], Bose–Chaudhuri–Hocquenghem (BCH) codes [48, 49] were used to construct efficient universal DD sequences for $k$-local Hamiltonians acting on physical qubits. By applying their construction to the quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ and assuming 1-local control Hamiltonians, we find that the length $L$ of a single DD cycle required to universally decouple $k$-local Hamiltonians scales as
@@
In Section V, we will present different constructions based on alternative classical codes that generates more time-efficient decoupling groups for 2-local and 3-local Hamiltonians. While the asymptotic scaling remains the same as in Eqs. (25) and (26), our constructions enable improved constant prefactors compared to those based on BCH codes for the 2-local and 3-local cases.
In the next section, we tighten the connection between DD and error detection codes by drawing an analogy to quantum stabilizer codes. With this viewpoint, if the structure of the noise model is known, or if the goal is to engineer a specific Hamiltonian, it becomes possible to design significantly more time-efficient DD sequences tailored to the problem.
@@ IV.3 Dynamical decoupling as error detection codes
Historically, self-dual additive codes over $\mathbb{F}_{4}$ (i.e., $\mathcal{C}=\mathcal{C}^{\perp}$) have played a central role in the construction of stabilizer quantum error-correcting codes [42, 47, 50, 51]. Many properties of stabilizer codes can be analyzed through their stabilizer generators. Similarly, in our setting, we will study the properties of the DD sequence through the generators of the associated additive classical code.
@@
Because $\mathcal{C}$ is an additive code, it is generated by a finite set of independent generators $\Gamma=\{\gamma_{1},\dots,\gamma_{m}\}$, where $|\Gamma|=\log_{2}|\mathcal{C}|$. Importantly, if a vector $\mathbf{u}\in\mathbb{F}_{4}^{n}$ anticommutes with at least one of the generators $\gamma_{i}$, then the Pauli string associated with $\mathbf{u}$ will be eliminated by the decoupling sequence generated from the code $\mathcal{C}$ [21].
This can be seen explicitly through a simple example. Suppose $\mathcal{C}$, and equivalently the DD group $\mathcal{G}$, has two generators $\gamma_{1}$ and $\gamma_{2}$. The DD group is then
$\{I,\gamma_{1},\gamma_{2},\gamma_{1}\gamma_{2}\}\ .$ (27)
@@
We have shown that DD sequences constructed from additive classical codes behave analogously to quantum stabilizer codes. However, we emphasize two key differences. First, unlike quantum stabilizer codes, we do not require the generators of the decoupling group $\mathcal{G}$ to commute with each other. Second, in the context of dynamical decoupling, there is no active "correction" step. Instead, the twirling procedure passively removes all components of the Hamiltonian that anticommute with at least one of the generators of $\mathcal{G}$. In this sense, the role of the decoupling sequence is similar to that of an error detection code rather than an error correction code.
The connection between designing decoupling sequences and designing error detection codes is a key insight of our work. Specifically, if the dominant residual interactions in the system are known, we can forego a universal decoupling group $\mathcal{G}$, which suppresses all interactions, and instead remove certain generators from $\mathcal{G}$ while still detecting the dominant residual terms [20]. The resulting DD sequence becomes significantly more time-efficient. We will make extensive use of this idea in Sections V.1 and V.2, where we tailor our general method to specific cases.
As an example, the orthogonal array presented in Fig. 4 that universally decouples all 2-local interactions is generated by the DD group
@@
However, if we know that the residual two-qubit interactions are restricted to terms of the form $X_{i}X_{j},Y_{i}Y_{j},$ or $Z_{i}Z_{j}$, as is the case e.g. in spin qubits [52, 53, 54, 55], the reduced decoupling group
@@
still detects all the relevant interactions. Explicitly, the elements of $\mathcal{G}_{\rm red}$ are
@@
We note that this group corresponds to running the standard $\rm XY4$ sequence in parallel on all three qubits, with the $\rm XY4$ cycles on the second and third qubits being cyclically shifted relative to each other. Moreover, the reduced group $\mathcal{G}_{\rm red}$ universally decouples all single-qubit terms. As a result, the DD sequence requires only $2^{2}=4$ time steps instead of $2^{4}=16$, achieving a fourfold reduction in cycle length.
Another interesting consequence of our error-detection perspective is that it enables the construction of decoupling sequences that selectively suppress certain interactions while preserving others [56, 24]. For instance, in systems where anisotropic exchange interactions arise, such as spin qubits with strong spin-orbit coupling [57, 56, 58], the decoupling sequence generated by Eq. (30) preserve specific components of the antisymmetric Dzyaloshinskii–Moriya (DM) interaction. In particular, it maintains terms of the form:
@@
By alternating between DD sequences derived from $\mathcal{G}_{\rm red}$ and suitably permuted variants of it, one can fully preserve the complete DM interaction, thus enabling new possibilities for quantum simulation [59, 60].
We will discuss more examples of sequences simulating certain Hamiltonians in Section V.3. In the next section, we introduce a search algorithm to construct selective DD sequences.
Finally, we remark that if one uses a quantum stabilizer code (i.e a self-dual additive code) to construct a DD sequence in parallel with employing the same code for active error correction, it is possible to improve the overall code threshold by suppressing high-weight errors. This idea was explored in Refs. [20, 21], where the authors designed decoupling sequences based on the stabilizers and logical operators of a quantum code to suppress high-weight error strings. We do not directly compare our sequences with those proposed in Refs. [21, 20], as they serve fundamentally different purposes. The sequences in that work scale exponentially with the quantum code (system) size but are used alongside a quantum error-correcting code, such that errors that accumulate due to the long sequence length can be actively corrected. In contrast, our sequences are designed to protect bare physical qubits without additional quantum error correction. An interesting direction for future work is combine our approach with that of Refs. [20, 21], to explore trade-offs between sequence length and robustness against errors in the presence of a quantum code.

### v2

@@ IV.1 Orthogonal arrays and dynamical decoupling
Time-efficient DD sequences can be constructed using combinatorial design objects. For example, Ref. [brown2024efficient] employs Hadamard matrices to construct decoupling sequences for $2$-local Hamiltonians. We generalize this approach to arbitrary $k$-local Hamiltonians by using more general combinatorial structures known as orthogonal arrays [Bookatz_2016, Rotteler_2006, Stollsteimer2001]. In the following analysis, we consider the worst-case scenario where all possible $k$-local interactions between the color classes $c_{i}$ are present in the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$.
@@
An orthogonal array $\text{OA}(L,\chi_{I},S,k)$ is an $L\times\chi_{I}$ array with entries taken from a finite set $S$ and strength $k$, such that every possible $k$-tuple from $S^{\otimes k}$ appears uniformly across any choice of $k$ columns [hedayat2012orthogonal]. In this work, we restrict ourselves to the set $S=\{I,X,Y,Z\}$ of single-qubit Pauli operators. Each column of the orthogonal array is associated with a color class $c_{i}$, and each row defines a group element in the decoupling group $\mathcal{G}$. If the noise model is biased along a specific axis of the Bloch sphere (e.g. the $Z$-axis) or if we want to maintain certain interactions (e.g. $XX$ interactions), we can reduce the required size $L$ of the orthogonal array by selecting a smaller set (e.g. $S=\{I,X\}$) instead of the full single-qubit Pauli group. We remark that for any choice of parameters $\chi_{I}$, $S$, and strength $k$, there always exists an orthogonal array whose length $L$ is determined by these parameters [hedayat2012orthogonal].
@@
In Fig. 4, we show the explicit orthogonal array $\text{OA}(L=16,\chi_{I}=3,|S|=4,k=2)$, which generates a universal decoupling sequence for 2-local Hamiltonians defined on the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$ [neilsloaneOrthogonalArrays]. One can verify that for any pair of columns, the rows of the OA exhaustively span all combinations of $\{I,X,Y,Z\}^{\otimes 2}$ exactly once. Moreover, each Pauli operator appears exactly four times in every column, ensuring balanced coverage and uniformity.
@@
By construction, the action of the decoupling group $\mathcal{G}$ derived from an OA with strength $k$ implements a group twirl over the $k$-fold Pauli group $\{I,X,Y,Z\}^{\otimes k}$ on any subset of $k$ color classes. As a consequence, all $k$-local interaction terms are dynamically suppressed, following directly from Schur’s lemma [Eq. (11)]. Furthermore, all interactions with support on less than $k$ color classes are also suppressed as a consequence of this construction. Therefore, to universally decouple all $k$-local interactions on the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$, it suffices to construct the orthogonal array $\text{OA}(L,\chi_{I},S,k)$.
@@
We note that there is a comprehensive online library cataloging efficient (i.e good ratio between $\chi_{I}$ and $L$) orthogonal arrays across a wide range of parameters [neilsloaneOrthogonalArrays]. In the next section, we describe how to construct length-efficient orthogonal arrays by utilizing their connection to classical error correction codes [Rotteler_2006, Bookatz_2016]. Since classical codes are extensively studied, we can utilize well-established classical codes to obtain time-efficient DD sequences suppressing arbitrary $k$-local interactions [Bookatz_2016].
@@ IV.2 Orthogonal array and classical additive codes
The connection between orthogonal arrays and classical linear codes is well established [DELSARTE1973407, hedayat2012orthogonal, Rotteler_2006, Bookatz_2016]. In this work, we focus on the projective Pauli group, which admits a natural isomorphism to the finite field $\mathbb{F}_{4}$. A finite field $\mathbb{F}_{q}$ is a set containing $q$ elements in which addition, subtraction, multiplication, and division (except by zero) are all well-defined and satisfy the field axioms. The isomorphism between the projective Pauli group and $\mathbb{F}_{4}$ is defined by the correspondence
@@
This correspondence allows us to represent Pauli operators algebraically as elements of $\mathbb{F}_{4}$, providing a convenient framework for constructing orthogonal arrays from quaternary codes. Furthermore, when the noise is biased (e.g. along the $Z$-axis), we may restrict to a subset of the Pauli group, in which case classical codes over the binary field $\mathbb{F}_{2}$ become relevant.
We consider here classical additive codes, which, as we will show in Sec. IV.3, provide a natural framework for interpreting a decoupling sequence as an error detection code. An additive code $\mathcal{C}$ is a subset of $\mathbb{F}_{q}^{n}$ (a vector of length $n$ in which its entries are from $\mathbb{F}_{q}$) that is closed under addition. While every linear code is an additive code, the converse is not true: additive codes are not necessarily closed under scalar multiplication unless the field is $\mathbb{F}_{2}$.
@@
The vector space $\mathbb{F}_{4}^{n}$ can be endowed with an inner product $\langle\cdot,\cdot\rangle$. A natural choice for additive codes is the trace Hermitian inner product $\langle\cdot,\cdot\rangle_{\rm tr}:\mathbb{F}_{4}^{n}\to\mathbb{F}_{2}$ [ezerman2011additive], defined as
$\langle\bf{u},\bf{v}\rangle_{tr}~=\sum_{i}(u_{i}v_{i}^{2}+v_{i}u_{i}^{2}).$ (19)
Under the mapping between single-qubit Pauli operators and elements of $\mathbb{F}_{4}$ given in Eq. (15), this inner product directly gives the commutation relations of Pauli operators. That is, it evaluates to zero if the corresponding Pauli operators commute and to one if they anticommute.
@@
$\mathcal{C}^{\perp}=\{\bf{v}\in\mathbb{F}_{q}^{n}~|~\langle\bf{u},\bf{v}\rangle_{tr}=0,\qquad\forall\bf{u}\in\mathcal{C}\}.$ (20)
@@
A seminal result by Delsarte [DELSARTE1973407] establishes a fundamental connection between orthogonal arrays and classical codes:
@@
In Ref. [Bookatz_2016], Bose–Chaudhuri–Hocquenghem (BCH) codes [bose1960class, gorenstein1961class] were used to construct efficient universal DD sequences for $k$-local Hamiltonians acting on physical qubits, not color classes. By applying their construction to the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$, we find that the length $L$ of a single DD cycle required to universally decouple $k$-local Hamiltonians scales as
@@
In Sec. V, we will present different constructions based on alternative classical codes that generate more time-efficient decoupling groups for 2-local and 3-local Hamiltonians. While the asymptotic scaling remains the same as in Eqs. (21) and (22), our constructions lead to improved constant prefactors compared to those based on BCH codes for the 2-local and 3-local Hamiltonians.
@@ IV.3 Dynamical decoupling as error detection codes
Historically, self-dual additive codes over $\mathbb{F}_{4}$ (i.e., $\mathcal{C}=\mathcal{C}^{\perp}$) have played a central role in the construction of stabilizer quantum error-correcting codes [calderbank1997quantumerrorcorrectioncodes, ezerman2011additive, zeng2011transversality, haah2016algebraic]. Many properties of stabilizer codes can be analyzed through their stabilizer generators. In a closely analogous manner, we further develop the connection between dynamical decoupling and classical additive codes by showing that key properties of a DD sequence can be analyzed through the generators of its decoupling group. This perspective not only provides a compact characterization of DD sequences, but also enables systematic constructions of selective DD sequences, as we explain in this section.
@@
Because $\mathcal{C}$ is an additive code, it is generated by a finite set of independent generators $\Gamma=\{\gamma_{1},\dots,\gamma_{m}\}$, where $|\Gamma|=\log_{2}|\mathcal{C}|$. Importantly, if a vector $\mathbf{u}\in\mathbb{F}_{4}^{n}$ anticommutes with at least one of the generators $\gamma_{i}$, then the Pauli string associated with $\mathbf{u}$ will be suppressed by the decoupling sequence generated from the code $\mathcal{C}$ [vezvaee2025demonstrationhighfidelityentangledlogical].
This can be seen explicitly through a simple example. Suppose $\mathcal{C}$, and equivalently the decoupling group $\mathcal{G}$, has two generators $\gamma_{1}$ and $\gamma_{2}$. The decoupling group is then given by
$\mathcal{G}=\{I,\gamma_{1},\gamma_{2},\gamma_{1}\gamma_{2}\}\ .$ (23)
@@
We have shown that DD sequences constructed from classical additive codes behave analogously to quantum stabilizer codes. However, we emphasize two key differences. First, unlike quantum stabilizer codes, we do not require the generators of the decoupling group $\mathcal{G}$ to commute with each other. Second, in the context of dynamical decoupling, there is no active "correction" step. Instead, the twirling procedure passively removes all components of the Hamiltonian that anticommute with at least one of the generators of $\mathcal{G}$. In this sense, the role of the decoupling sequence is similar to that of an error detection code rather than an error correction code.
The connection between designing decoupling sequences and designing error detection codes is a key insight of our work. Specifically, if the dominant residual interactions in the system are known, we can forego a universal decoupling group $\mathcal{G}$, which suppresses all interactions, and instead remove certain generators from $\mathcal{G}$ while still suppressing the dominant residual terms [paz2013optimally]. The resulting DD sequence length becomes significantly shorter. We will make extensive use of this idea in Sec. V.1 and V.2, where we tailor our general method to specific cases.
As an example, the orthogonal array presented in Fig. 4 that universally decouples all 2-local interactions is generated by the decoupling group
@@
However, if we know that the residual two-qubit interactions are restricted to terms of the form $X_{i}X_{j},Y_{i}Y_{j},$ or $Z_{i}Z_{j}$, as in spin qubits [burkard2023semiconductor, Xue2022, noiri2022fast, mills2022two], the reduced decoupling group
@@
still suppress all the relevant residual interactions. Explicitly, the elements of $\mathcal{G}_{\rm red}$ are
@@
We note that this group corresponds to running the standard $\rm XY4$ sequence in parallel on all three qubits, with the $\rm XY4$ cycles on the second and third qubits being cyclically shifted relative to each other. Moreover, the reduced group $\mathcal{G}_{\rm red}$ still universally decouples all single-qubit terms. As a result, the DD sequence requires only $2^{2}=4$ time steps instead of $2^{4}=16$, achieving a fourfold reduction in cycle length.
Another interesting consequence of our error-detection perspective is that it enables the construction of decoupling sequences that selectively suppress certain interactions while preserving others [nguyen2025single, evert2024syncopated]. For instance, in systems where anisotropic exchange interactions arise, such as spin qubits with strong spin-orbit coupling [Geyer2024, nguyen2025single, saezmollejo2024exchangeanisotropiesmicrowavedrivensinglettriplet], the decoupling sequence generated by Eq. (26) preserves specific components of the antisymmetric Dzyaloshinskii–Moriya (DM) interaction. In particular, it maintains terms of the form:
@@
By alternating between DD sequences derived from $\mathcal{G}_{\rm red}$ and suitably permuted variants of it, one can fully preserve the complete DM interaction, thus enabling new possibilities for quantum simulation [fariña2025siteresolvedmagnontriplondynamics, Hsiao].
We will discuss more examples of sequences simulating certain Hamiltonians in Sec. V.3. In the next section, we introduce a simple search algorithm to construct selective DD sequences.
Finally, we remark that if one uses a quantum stabilizer code (i.e a self-dual additive code) to construct a DD sequence in parallel with employing the same code for active quantum error correction, it is possible to improve the overall quantum code performance by suppressing high-weight errors. This idea was explored in Refs. [paz2013optimally, vezvaee2025demonstrationhighfidelityentangledlogical], where the authors designed decoupling sequences based on the stabilizers and logical operators of a quantum code to suppress high-weight error strings. We do not directly compare our sequences with those proposed in Refs. [vezvaee2025demonstrationhighfidelityentangledlogical, paz2013optimally], as they serve fundamentally different purposes. The sequences in their work scale exponentially with the quantum code (system) size but are used alongside a quantum error-correcting code, such that errors that accumulate due to the long sequence length can be actively corrected. In contrast, our sequences are designed to protect bare physical qubits without additional quantum error correction. An interesting direction for future work is combining our approach with that of Refs. [paz2013optimally, vezvaee2025demonstrationhighfidelityentangledlogical], to explore trade-offs between sequence length and robustness against errors in the presence of a quantum stabilizer code.